OpenStudy (anonymous):

If f(x) = 3x + 1 and g(x) = 2x − 4, find f[g(2)]. 25 4 0 1

OpenStudy (anonymous):

first find g(2)

OpenStudy (anonymous):

then the value which comes substitute it

OpenStudy (anonymous):

g(2)=2(2)-4 and f[g(2])= 3[g(2)] + 1

OpenStudy (anonymous):

im still confused

OpenStudy (anonymous):

g(x) is your inside function, so you plug in 2 there

OpenStudy (anonymous):

what is the value of g(2) @isaiahw.

OpenStudy (anonymous):

4f?

OpenStudy (anonymous):

g(2) = 2x2-4 = 0

OpenStudy (anonymous):

thus f(g[2]) = f(0) = 3x0 + 1 = 1

OpenStudy (anonymous):

First get g(2) from g(x) given then you get a value.. put that value in f(x) and calculate